8. Assume a single firm in a purely competitive industry has costs as indicated in the following table.填表TFC: 40 TP TVC 正兼利3 肇煳サ iminat AVC ATC 40 N/A N/A N/A N/A 75 90 110 20 3 1 1332 3 洋性 170 2lo hbb 2833 220 290 3205 (a) At a product price of S65, will this firm produce in the short run? What wil its profit or loss be? What will (b) At a product price of $18, will this firm produce in the short run? its profit or loss be?
Answer #1

Answer - 8

ATC = TC/Q , AVC = TVC/Q , MC = Delta TC/Delta Q

TC = TVC + TFC,

Using these formulas we have the following result:

TP TVC TC AFC AVC ATC MC 95 57.5 30 43.33333 37.5 95 115 37.5 2 130 13.33333 110 135 170 27.5 27 210 6.666667 28.33333 10 175 260 5.714286 31.42857 37.14286 5 330 36.25 41.25

a)

A purely competitive firm produce that quantity at which Price(P) = MC( and If P not equals MC then P must be just greater than MC) and MC is rising. Also in the short run A firm will prefer to shut down If P is lesser than AVC.

Suppose P = 65, then P will be just greater than MC and MC is rising when Q = 7.

Hence Firm will produce 7 units.

Profit = Total Revenue - Total Cost = P*Q - TC = 7*65 - 260 = $195

b)

Suppose P = 18, then P will be just greater than MC , Q = 3. Note that when Q = 3 , AVC = 30

Hence, AVC > P. This implies firm will prefer to shutdown.

Hence Firm will produce 0 units.

Profit = Total Revenue - Total Cost = P*Q - TC = 3*0 - 40 = -$40

Hence. Firm's loss = $40 (or Profit = $-40)